CS294-3: Algorithms in the real world

Name: Email: Homepage: Department (division): Advisor (if relevant): Credit or Audit? Would you rather scribe one class, or give a 35 minute lecture on one of the topics? Which of the following algorithms/topics have you studied before

1. Fast Fourier Transform

2. Minimum-spanning-trees

3. Strassen’s matrix multiply

4. Factoring algorithms

5. RSA

6. Lempel-Ziv Compression

7. Simplex method

8. Computational geometry

9. Nbody algorithms